OpenStudy (anonymous):

Choose the equation of the line passing through the point (2, -4) and parallel to y = 3x - 6. y = 3x - 10 y = 3x - 6 y = 3x - 2 y = 3x - 12

OpenStudy (zzr0ck3r):

plug in the values and see if it works y = -4 = 3(2)-10

OpenStudy (anonymous):

If it's parallel then the slopes are the SAME but the y-intercepts are DIFFERENT So you know the slope in the new equation will be 3.

OpenStudy (anonymous):

Which is obvious. Now use point-slope form or do what @zzr0ck3r said.

OpenStudy (anonymous):

Point-Slope Form: \[y-y _{1}=m(x-x _{1})\]

OpenStudy (anonymous):

Where (2, -4) are your \[(x _{1},y _{1})\]

OpenStudy (anonymous):

Is it D?

OpenStudy (anonymous):

B

OpenStudy (anonymous):

O

OpenStudy (anonymous):

ok